# Arrows of Time for Large Language Models

### v2

@@ Abstract
We study the probabilistic modeling performed by Autoregressive Large Language Models through the angle of time directionality. We empirically find a time asymmetry exhibited by such models in their ability to model natural language: a difference in the average log-perplexity when trying to predict the next token versus when trying to predict the previous one. This difference is at the same time subtle and very consistent across various modalities (language, model size, training time, …). Theoretically, this is surprising: from an information-theoretic point of view, there should be no such difference. We provide a theoretical framework to explain how such an asymmetry can appear from sparsity and computational complexity considerations, and outline a number of perspectives opened by our results.
@@ 1.1 Autoregressive LLMs
Autoregressive LLMs (GPTs, GRUs, LSTMs, …) thus factorize (their estimates of) the joint probabilities in terms of conditional probabilities for each token knowing past ones. This brings a number of advantages: first, leverage the fact that for each token sequence $x_{1},\ldots,x_{n}$, each token $x_{k}$ is used in to predict each token $x_{\ell}$ with $\ell>k$. In particular, GPT (compared to the earlier BERT) includes causality-aware attention, allowing for a parallelization of the training process: a sequence $x_{1},\ldots,x_{n}$ generates $n-1$ fully parallelizable tasks (predict $x_{k}$ from $\left(x_{1},\ldots,x_{k-1}\right)$ for $2\leq k\leq n$). Also, this representation enables a natural sampling from $\mathbb{P}_{n}^{\rightarrow}$ (token by token), as well as data compression: the factorization decomposes these processes into many smaller substeps (Graves et al., 2023).
@@ 1.3 Cross-Entropy Loss and Perplexity
It is worth noting that, in spite of its apparent triviality, Remark 2 crucially depends on the choice $\ell$ as $\ell_{\mathcal{C}}$. Moreover, even if $\mathbb{P}_{n}^{\rightarrow}=\mathbb{P}_{n}^{\leftarrow}$, we will generally have $\ell_{i}^{\rightarrow}\neq\ell_{i}^{\leftarrow}$ for $1\leq i\leq n$ (though $\sum_{i}\ell_{i}^{\rightarrow}=\sum_{i}\ell_{i}^{\leftarrow}$). When $\mathbb{P}_{n}^{\rightarrow}=\mathbb{P}_{n}^{\leftarrow}=\mathbb{P}_{n}$ we have that $\ell_{i}^{\rightarrow}$ and $\ell_{i}^{\leftarrow}$ yield two (typically different) decompositions of the log-likelihood of $\left(x_{1},\ldots,x_{n}\right)$. For instance, take as a dataset the 81 expressions $A\times B=CD$ for $1\leq A,B\leq 9$ (setting $C=0$ when needed). The FW log-perplexity is concentrated on $A$ and $B$, each contributing $\ln{9}\approx 2.2$ nats: $\ell_{A}^{\rightarrow}=\ell_{B}^{\rightarrow}\approx 2.2$, $\ell_{C}^{\rightarrow}=\ell_{D}^{\rightarrow}=0$. The BW log-perplexity is distributed differently: for instance, for $3\times 4=12$, $(\ell_{A}^{\leftarrow},\ell_{B}^{\leftarrow},\ell_{C}^{\leftarrow},\ell_{D}^{\leftarrow})\approx(0,1.39,1.1,1.91)$.
@@ 1.4 Setup and Plan
1.5 Relation to Previous Works
@@
The idea of training language models ‘backward’, on reverse sequences, has occurred in a number of contexts. A notable example is (Sutskever et al., 2014), focusing on machine translation using LSTMs, finding that reversing the source sentence (i.e. the source LSTM learning backward) greatly improves performance. Other well-known examples of related techniques include ULMFiT (Howard & Ruder, 2018) and ELMO (Peters et al., 2018).
@@
Attempts at combining FW/BW models include (Mou et al., 2016; Liu et al., 2016; Zhang et al., 2018; Serdyuk et al., 2018; Mangal et al., 2019) (using RNNs); or recently (Nguyen et al., 2023), a ‘Meet in the Middle’ approach which shows how pre-training using FW/BW Transformer models enhance FW-only autoregressive generation; as well as in (Shen et al., 2023), applying the idea of reversing data to improve LLM performance (see also Section 3.1). In these approaches, the FW/BW models are treated as one model, yielding one combined loss. They compare various models on a task (see section 5 of (Nguyen et al., 2023) for a comprehensive review), rather than study potential discrepancies in FW/BW learning, as in most papers combining FW and BW generation.
A recent paper on adversarial attacks on LLMs (Pfau et al., 2023) compares FW/BW LLMs (with a BW GPT, 160M parameters), and an older study (Duchateau et al., 2002), uses BW N-gram language models for continuous speech recognition (comparing FW/BW trigrams models). In those, the authors report that BW models perform respectively equally or slightly better than FW models, in contradiction with our main findings. This does not affect our confidence in our results, given the magnitude of our experiments and the level of care involved in their setup.
@@ 2.1 Setup
For the identification of an AoT in a dataset, we make sure that both the FW and BW models are trained with the exact same specifications: the only reason for a difference between the models’ performances is hence the token prediction order. In all experiments, the models are trained from scratch, using He initialization.
@@ 2.1.2 Models, Hyperparameters and Training
While some experiments involve other autoregressive models (GRU, LSTM), for most training jobs we use the decoder-only Transformer (GPT) (Radford et al., 2018); our implementation (with code in the supplementary material) is derived from minGPT (Karpathy, 2020). All GPT experiments use learned positional embeddings and a dropout rate of $0.1$. Other hyperparameters may depend on the experiment, see the Appendix.
For all models, we use the AdamW optimizer (Loshchilov & Hutter, 2019) with base learning rate of $10^{-4}$ and a learning rate schedule with a warmup, followed by cosine annealing with warm restarts (Loshchilov & Hutter, 2017). These hyperparameters are mostly kept constant across different experiments, although the period of the warm restarts might be tweaked to synchronize the end of training with the end of a cycle, see the Appendix.
@@ 2.2.1 Arrow of Time in English and French
We begin in this section by analyzing the difference in FW vs BW training dynamics for a Transformer of size GPT2-Medium ($\sim 405M$ parameters, context window length of $256$) on the CC100 datasets for English and French. We train the FW and BW models for the equivalent of 1 epoch of the French dataset ($\sim 27B$ tokens), avoiding memorization.
@@ 2.2.5 Other Languages
From Table 3, we can see that in all the cases we tested, a FW AoT emerges, although its magnitude appears to vary from language to language, suggesting some universality of this phenomenon across human languages.
2.2.6 Possible Tokenization Artifacts
While the training procedures are perfectly symmetric with respect to the two directions, it is important to rule out any other possible sources of asymmetry. One possible source could in principle be the tokenization; indeed, the BPE tokenizer is trained in the FW direction. To rule out this possibility, we inverted (at the character level) one dataset (Greek), and re-trained a BPE tokenizer on the result. We trained a GPT2-Medium on it, and confirmed that the direction of the BPE tokenization has no effect on the training dynamics: in this case, the FW (respectively BW) model performs exactly the same way as the BW (respectively FW) model on the original tokenization, thereby showing exactly the same AoT. See Appendix A for details.
@@ 3.1.1 Synthetic dataset
For fixed $k\geq 1$, consider the synthetic language of strings of the form $p\times q\leftrightarrow\mathrm{rev}\left(pq\right)$, with $p<q$ primes, $p,q<10^{k}$ and $\mathrm{rev}\left(pq\right)$ being the product of $p$ times $q$, written in reverse order (see (Shen et al., 2023; Lee et al., 2023)). The numbers $p$ and $q$ are padded to be of $k$ digits exactly and the $\mathrm{rev}\left(pq\right)$ is padded to be of $2k$ digits exactly (e.g. for $k=4$: $0019\times 0023\leftrightarrow 73400000$). The symbols $\times$ and $\leftrightarrow$ are written as multiple tokens (3 tokens for $\times$, and 7 tokens for $\leftrightarrow$), to facilitate the learning of non-trivial operations by GPTs. For a fixed $k$, $\mathbb{P}$ is thus supported on sequences of $104k+$ 10 tokens; in our experiments, we set $k=5$ and take a dataset containing $10^{8}$ such random ordered pairs. Intuitively, computing the RHS of the symbol $\leftrightarrow$ given the LHS should be relatively easy (or at least learnable), while computing the LHS from the RHS should be much harder (at least finding the factor $q$; given $\mathrm{rev}\left(pq\right)$ and $q$, $p$ should be relatively easy to find).
@@ 3.1.4 Discussion
The above setup shows a significant AoT for the $p\times q\leftrightarrow\mathrm{rev}\left(pq\right)$ dataset. This discrepancy can be largely attributed to the asymmetry between the difficulty of factoring versus multiplication: compared to the information-theoretical limit, about $4.55$ nats are lost for the multiplication, while $8.43$ nats are lost for the factorization. We also see that the different structures of the LHS and RHS (which have the same information-theoretic content as they determine each other) also present a significant difference w.r.t. the models’ abilities: while the FW model reaches essentially optimal perplexity for the LHS (i.e. it recognizes primes $<10^{k}$), the BW model is very far from optimal on the RHS (i.e. to recognize products of primes pairs $p,q<10^{k}$ proves to be much more difficult). While part of the asymmetry is attributable to the models’ specifics, as long as the dataset size is kept so high that all pairs $\left(p,q\right)$ cannot be memorized, a significant AoT can be expected: as the model size (and training time) grows, multiplication will eventually be learned (long before the dataset can be memorized, (Shen et al., 2023)), while extracting substantial information from $\mathrm{rev}\left(pq\right)$ about $q$ should remain very hard. The above data model displays an AoT of types (1) and (2) (see 3 above): certain features turn out to be harder to learn for the BW models, while others simply are likely not even representable by LLMs (of reasonable sizes), as the alternative would yield an efficient factorization algorithm.
@@ 3.2 Binary Operations
The model of Section 3.1 shows how an AoT can appear in a dataset: in that example, based on computational complexity ideas, we could handcraft a synthetic dataset that is both practically and theoretically harder for a BW model than a FW one. This still leaves the question of why an AoT would arise in a dataset such as natural language, as in Section 2, and why in one direction rather than another, i.e. why FW models would consistently outperform BW ones. In this section, we introduce synthetic datasets based on operations on the space of $m$-bit registers identified with $\mathbb{F}_{2}^{m}$ ($\mathbb{F}_{2}$ denotes the field of integers mod $2$). We focus on languages based on $\mathbb{F}_{2}$-linear circuits and relate their learnability to their sparsity and use this to provide an explanation for a difference between FW and BW learnability; we then provide a setup motivating the specific FW direction observed on the AoT in natural languages; we then provide experiments validating our framework; we conclude by discussing extensions to the nonlinear case.
@@ 3.2.1 Linear Sparse Circuit Dynamics
Consider the class of measures $\mathbb{P}_{n}$ on a linear language formed by sequences of $n=2m+1$ tokens, of the form $x\leftrightarrow y$ where $x,y$ are random uniform on $\mathbb{F}_{2}^{m}$, but related by a bijective linear map; $\leftrightarrow$ is counted as a token. For each $\mathbb{P}_{n}$, we can write $y=f_{\mathbb{P}_{n}}^{\rightarrow}\left(x\right)$ and $x=f_{\mathbb{P}_{n}}^{\leftarrow}\left(y\right)$ for $f^{\leftrightarrows}:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$. We define the sparsity of a linear map $f:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$ as the proportion of zero entries of its matrix. We will (informally) say that a matrix is sparse if this proportion is relatively high, i.e. close to $1$. Intuitively, the sparsity of $f^{\rightarrow}$ (resp. of $f^{\leftarrow}$) is related to how easy it is to learn $\mathbb{P}_{n}$ (based on random data samples) for a FW model (resp. for a BW model). For GPT predictors, this is studied numerically in Section 3.2.3 below. For a linear language $\mathbb{P}_{n}$, an AoT will thus emerge if the sparsities of $f^{\rightarrow}$ and $f^{\leftarrow}$ are significantly different. It is common knowledge that the inverse of a sparse matrix is generally less sparse (see e.g. (Duff et al., 2017), Section 15.6), This is the basis for the following claim (verified numerically in Appendix B.2.2):

### v3

@@ Abstract
We study the probabilistic modeling performed by Autoregressive Large Language Models (LLMs) through the angle of time directionality, addressing a question first raised in (Shannon, 1951). For large enough models, we empirically find a time asymmetry in their ability to learn natural language: a difference in the average log-perplexity when trying to predict the next token versus when trying to predict the previous one. This difference is at the same time subtle and very consistent across various modalities (language, model size, training time, …). Theoretically, this is surprising: from an information-theoretic point of view, there should be no such difference. We provide a theoretical framework to explain how such an asymmetry can appear from sparsity and computational complexity considerations, and outline a number of perspectives opened by our results.
@@ 1.1 Autoregressive LLMs
Autoregressive LLMs (GPTs, GRUs, LSTMs, …) thus factorize (their estimates of) the joint probabilities in terms of conditional probabilities for each token knowing past ones. This brings a number of advantages: first, this leverages the fact that for each token sequence $x_{1},\ldots,x_{n}$, each token $x_{k}$ is used in to predict each token $x_{\ell}$ with $\ell>k$. In particular, GPT (compared to the earlier BERT) includes causality-aware attention, allowing for a parallelization of the training process: a sequence $x_{1},\ldots,x_{n}$ generates $n-1$ fully parallelizable tasks (predict $x_{k}$ from $\left(x_{1},\ldots,x_{k-1}\right)$ for $2\leq k\leq n$). Also, this representation enables a natural sampling from $\mathbb{P}_{n}^{\rightarrow}$ (token by token), as well as data compression: the factorization decomposes these processes into many smaller substeps (Graves et al., 2023).
@@ 1.3 Cross-Entropy Loss and Perplexity
It is worth noting that, in spite of its apparent triviality, Remark 2 crucially depends on the choice $\ell$ as $\ell_{\mathcal{C}}$. Moreover, even if $\mathbb{P}_{n}^{\rightarrow}=\mathbb{P}_{n}^{\leftarrow}$, we will generally have $\ell_{i}^{\rightarrow}\neq\ell_{i}^{\leftarrow}$ for $1\leq i\leq n$ (though $\sum_{i}\ell_{i}^{\rightarrow}=\sum_{i}\ell_{i}^{\leftarrow}$). When $\mathbb{P}_{n}^{\rightarrow}=\mathbb{P}_{n}^{\leftarrow}=\mathbb{P}_{n}$ we have that $\ell_{i}^{\rightarrow}$ and $\ell_{i}^{\leftarrow}$ yield two (typically different) decompositions of the log-likelihood of $\left(x_{1},\ldots,x_{n}\right)$. For instance, take as a dataset the 81 expressions $A\times B=CD$ for $1\leq A,B\leq 9$, $0\leq C,D\leq 9$ (setting $C=0$ when needed). The FW log-perplexity is concentrated on $A$ and $B$, each contributing $\ln{9}\approx 2.2$ nats: $\ell_{A}^{\rightarrow}=\ell_{B}^{\rightarrow}\approx 2.2$, $\ell_{C}^{\rightarrow}=\ell_{D}^{\rightarrow}=0$. The BW log-perplexity is distributed differently: for instance, for $3\times 4=12$, $(\ell_{A}^{\leftarrow},\ell_{B}^{\leftarrow},\ell_{C}^{\leftarrow},\ell_{D}^{\leftarrow})\approx(0,1.39,1.1,1.91)$.
@@ 1.4 Setup and Plan
1.5 Relation to Previous Works in Language Modeling
To the best of our knowledge, the question of comparing FW and BW text generation in Language Modeling was first raised in (Shannon, 1951): Shannon ran experiments on the task of predicting the next vs previous letters, noting the theoretical equality between FW and BW entropies; he noted that while the BW prediction appeared to be “subjectively much more difficult” for humans, it led to “only slightly poorer” scores.
@@ 1.5 Relation to Previous Works in Language Modeling
A notable recent example is (Sutskever et al., 2014), focusing on machine translation using LSTMs, finding that reversing the source sentence (i.e. training the source LSTM backwards) improves performance. Other well-known examples of related techniques include ULMFiT (Howard & Ruder, 2018) and ELMO (Peters et al., 2018), the already mentioned BERT (Devlin et al., 2019), as well as T5 (Raffel et al., 2023), and XLNET (Yang et al., 2020).
@@
Attempts at combining FW/BW models include (Mou et al., 2016; Liu et al., 2016; Zhang et al., 2018; Serdyuk et al., 2018; Mangal et al., 2019) (using RNNs); or recently (Nguyen et al., 2023), a ‘Meet in the Middle’ approach which shows how pre-training using FW/BW Transformer models enhance FW-only autoregressive generation; as well as in (Shen et al., 2023), applying the idea of reversing data to improve LLM performance (see also Section 3.1). In these approaches, the FW/BW models are treated as one model, yielding one combined loss. They compare various models on a task (see section 5 of (Nguyen et al., 2023) for a comprehensive review), rather than study potential discrepancies in FW/BW learning.
Some results showing BW models performing equally or even better than FW models can be found in the literature. While (Vinyals et al., 2016) highlights the importance of order (of input and output sequences) for performance, and shows that scrambling words reduces performance, it shows FW and BW seemingly performing equally well. In a recent work (Pfau et al., 2023) use BW GPT models to perform adversarial attacks on LLMs, showing slightly better accuracy BW than FW. An older study (Duchateau et al., 2002) based on trigram models, also seemingly reports better BW than FW performances. Note that these works do not affect our confidence in our results, given the magnitude of our experiments and the level of care involved in their setup.
A number of works, in particularly related to the machine-translation setup, try to use token re-ordering to improve performances in one way or another, see, e.g. (Wu et al., 2018; Oord et al., 2017; Gu et al., 2018; Lee et al., 2018; Ford et al., 2018; Savinov et al., 2022; Welleck et al., 2019; Stern et al., 2019; Gu et al., 2019b; Chan et al., 2019b, a; Gu et al., 2019a; Emelianenko et al., 2019; Mansimov et al., 2020). See (Xiao et al., 2023) for a survey. Given the extensive body of work on the topic, it comes across as somehow surprising that the effect we highlight in our paper is not noted anywhere (besides in the early works of Shannon). Among possible reasons for this could be the use of translation-specific metrics such as the BLEU score (Papineni et al., 2001), rather than cross-entropy losses, and the lack of careful setups comparing FW and BW performance for large models on large datasets, all else being equal.
1.6 Causality and Information Theory
While the AoT effect we highlight in this paper is surprising from the point of view of information theory, there are several theoretical frameworks that appear to be related to this effect:
  • Structural Causal Models (Peters et al., 2017) consist of families of random variables linked by certain relationships that (implicitly) involve a notion of time: consider random variables $X_{1},\ldots,X_{n}$ such that for each $i\geq 1$, $X_{i+1}$ is a written as $f_{i}(X_{1},\ldots,X_{i},Z_{i})$, where the $Z_{i}$’s are jointly independent. While the presence of such a decomposition is not special to the order in which the variables are labeled, an order may be singled out in certain cases if we put constraints on the structure of $f_{i}$ and $Z_{i}$ (e.g. that $f_{i},Z_{i}$ are ‘simple’ in some sense): if we e.g. reverse the order of the variables $(\tilde{X}_{1},\ldots,\tilde{X}_{n})=(X_{n},\ldots,X_{1})$, it may not be possible to write $\tilde{X}_{i+1}=\tilde{f}_{i}(\tilde{X}_{1},\ldots,\tilde{X}_{i},\tilde{Z}_{i})$, with $\tilde{f}_{i},\tilde{Z}_{i}$ being as ‘simple’ as $f_{i},Z_{i}$. This may have an impact on learnability (Scholkopf et al., 2021), akin to that of an AoT.
  • Computationally-constrained views on information theory, in particular the recent framework of $\mathcal{V}$-information (Xu et al., 2020), allow one to take into account the computational challenges associated with the information extraction; through the lens of the $\mathcal{V}$-information, we see the emergence of symmetry breaking, an example of which is the Arrow-of-Time effect.
@@ 2.1 Setup
For the identification of an AoT in a dataset, we make sure that both the FW and BW models are trained with the exact same specifications: the only reason for a difference between the models’ performances is hence the token prediction order. In all experiments, the models are trained from scratch, using He initialization (He et al., 2015).
@@ 2.1.2 Models, Hyperparameters and Training
While some experiments involve other autoregressive models (GRU, LSTM), for most training jobs we use the decoder-only Transformer (GPT) (Radford et al., 2018); our implementation (with code in the supplementary material) is derived from minGPT (Karpathy, 2023). All GPT experiments use learned positional embeddings and a dropout rate of $0.1$. Other hyperparameters may depend on the experiment, see the Appendix.
For all models, we use the AdamW optimizer (Loshchilov & Hutter, 2019) with base learning rate of $10^{-4}$ and a learning rate schedule with a warmup, followed by cosine annealing with warm restarts (Loshchilov & Hutter, 2017). These hyperparameters are mostly kept constant across different experiments, although the period of the warm restarts might be tweaked to synchronize the end of training with the end of a cycle, see Appendix A for details.
@@ 2.2.1 Arrow of Time in English and French
We begin in this section by analyzing the difference in FW vs BW training dynamics for a Transformer of size GPT2-Medium ($\sim 405M$ parameters, context window length of $256$) on the CC-100 datasets for English and French. We train the FW and BW models for the equivalent of 1 epoch of the French dataset ($\sim 27B$ tokens), avoiding memorization.
@@ 2.2.5 Other Languages
Figure 3: Validation loss curves for FW and BW models during training. Consistently, the BW loss is higher than its FW counterpart. This persists through the warm restart of the learning rate, which causes a bump in the loss.
@@
From Table 3, we can see that in all the cases we tested, a FW AoT emerges, although its magnitude appears to vary from language to language, suggesting some universality of this phenomenon across human languages. Fig. 3 showcases the stability of this AoT during training, across languages. Three more languages (Tagalog, Hebrew and Arab) were tested at the suggestion of reviewers, confirming the universality of the AoT in human languages (see Appendix A.5).
2.2.6 Possible Artifacts
While the training procedures are perfectly symmetric with respect to the two directions, it is important to rule out any other possible sources of asymmetry. One possible source could in principle be the tokenization; indeed, the BPE tokenizer is trained in the FW direction. To rule out this possibility, we inverted (at character level) two datasets (Greek and French), and re-trained a BPE tokenizer on the result. We trained a GPT2-Medium on it, and confirmed that the direction of the BPE tokenization has no effect on the training dynamics: in this case, the FW (respectively BW) model performs very closely to the BW (respectively FW) model on the original tokenization, thereby showing exactly the same AoT. See Appendix A.6 for details.
Additionally, one might ask about the variation in the final losses due to initialization. Although the agreement of all the different experiments show that this is negligible w.r.t. the AoT effect, we quantify this influence by repeating experiments for Greek, see app. A.7.
@@ 3.1.1 Synthetic dataset
For fixed $k\geq 1$, consider the language of strings of the form $p\times q\leftrightarrow\mathrm{rev}\left(pq\right)$, with $p<q$ primes, $p,q<10^{k}$ and $\mathrm{rev}\left(pq\right)$ being the product of $p$ times $q$, written in reverse order (see (Shen et al., 2023; Lee et al., 2023)). The numbers $p$ and $q$ are padded to be of $k$ digits exactly and the $\mathrm{rev}\left(pq\right)$ is padded to be of $2k$ digits exactly (e.g. for $k=4$: $0019\times 0023\leftrightarrow 73400000$). The symbols $\times$ and $\leftrightarrow$ are written as multiple tokens (3 tokens for $\times$, and 7 tokens for $\leftrightarrow$), to facilitate the learning of non-trivial operations by GPTs (Thomas Ahle, 2023). For a fixed $k$, $\mathbb{P}$ is thus supported on $104k+$ 10 token sequences; in our experiments, we set $k=5$ and take $10^{8}$ such random ordered pairs. Intuitively, computing the right-hand side (RHS) of the symbol $\leftrightarrow$ given the left-hand side (LHS) should be easy, while computing the LHS from the RHS should be much harder (at least finding $q$; given $q$ and $\mathrm{rev}\left(pq\right)$, finding $p$ should be easier).
@@ 3.1.4 Discussion
The above setup shows a significant AoT for the $p\times q\leftrightarrow\mathrm{rev}\left(pq\right)$ dataset. This discrepancy can be largely attributed to the asymmetry between the difficulty of factoring versus multiplication: compared to the information-theoretical limit, about $4.55$ nats are lost for the multiplication, while $8.43$ nats are lost for the factorization. We also see that the different structures of the LHS and RHS (which have the same information-theoretic content as they determine each other) also present a significant difference w.r.t. the models’ abilities: while the FW model reaches essentially optimal perplexity for the LHS (i.e. it recognizes primes $<10^{k}$), the BW model is very far from optimal on the RHS (i.e. to recognize products of primes pairs $p,q<10^{k}$ proves to be much more difficult). While part of the asymmetry is attributable to the models’ specifics, as long as the dataset size is kept high enough that all pairs $\left(p,q\right)$ cannot be memorized, a significant AoT can be expected: as the model size (and training time) grows, multiplication will eventually be learned (long before the dataset can be memorized, (Shen et al., 2023)), while extracting substantial information from $\mathrm{rev}\left(pq\right)$ about $q$ should remain very hard. The above data model displays an AoT of types (1) and (2) (see 3 above): certain features turn out to be harder to learn for the BW models, while others simply are likely not even representable by LLMs (of reasonable sizes), as the alternative would yield an efficient factorization algorithm.
@@ 3.2 Binary Operations
The model of Section 3.1 shows how an AoT can appear in a dataset: in that example, based on computational complexity ideas, we could handcraft a synthetic dataset that is both practically and theoretically harder for a BW model than a FW one. This still leaves the question of why an AoT would arise in a dataset such as natural language, as in Section 2, and why in one direction rather than another, i.e. why FW models would consistently outperform BW ones. In this section, we introduce synthetic datasets based on operations on the space of $m$-bit registers identified with $\mathbb{F}_{2}^{m}$ ($\mathbb{F}_{2}$ denotes the field of integers mod $2$). We focus on languages based on $\mathbb{F}_{2}$-linear circuits and relate their learnability to their sparsity, using this to explain a difference between FW and BW learnability; we then provide a setup motivating the specific FW direction of the AoT in natural languages; we then provide experiments validating our framework; we conclude by discussing extensions to the nonlinear case.
@@ 3.2.1 Linear Sparse Circuit Dynamics
Consider the class of measures $\mathbb{P}_{n}$ on a linear language formed by sequences of $n=2m+1$ tokens, of the form $x\leftrightarrow y$ where $x,y$ are random uniform on $\mathbb{F}_{2}^{m}$, but related by a bijective linear map; $\leftrightarrow$ is counted as a token. For each $\mathbb{P}_{n}$, we can write $y=f_{\mathbb{P}_{n}}^{\rightarrow}\left(x\right)$ and $x=f_{\mathbb{P}_{n}}^{\leftarrow}\left(y\right)$ for $f^{\leftrightarrows}:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$. We define the sparsity of a linear map $f:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$ as the proportion of zero entries of its matrix. We will (informally) say that a matrix is sparse if this proportion is relatively high, i.e. close to $1$. Intuitively, the sparsity of $f^{\rightarrow}$ (resp. of $f^{\leftarrow}$) is related to how easy it is to learn $\mathbb{P}_{n}$ (based on random data samples) for a FW model (resp. for a BW model). For GPT predictors, this is studied numerically in Section 3.2.3 below. For a linear language $\mathbb{P}_{n}$, an AoT will thus emerge if the sparsities of $f^{\rightarrow}$ and $f^{\leftarrow}$ are significantly different. It is common knowledge that the inverse of a sparse matrix is generally less sparse (e.g. (Duff et al., 2017), Section 15.6). This is the basis for the following claim (verified numerically in Appendix B.2.2):